Question7 Billie travels 3.2 km due east in 0.1 hr, then 3.2 km at 15.0 degrees eastward of due north in 0.21 hr, and finally another 3.2 km due east in 0.1 hr. What is the average velocity for the entire trip? Select the correct answer CHECK ANSWER 0 of 1 attempts used LAST ATTEMPT O 7 km/hr O 3 km/hr O 13 km/hr O 19 km/hr O 25 km/hr
Answer #1

Let take east direction as +x axis and north direction as +y axis.

Therefore

rm.

S3.2km 15 degrees east of north

thus

S2r-3.2sin(15°)0.83km

and

S2.,-3·2cos(15°) = 3.09km

= 0.21hr

Finally

S3r3.2km, S 3y0,130.1hn

Now,

the total displacement in the x direction is

S_x = S_1_x +S_2_x +S_3_x and   S_y = S_1_y +S_2_y +S_3_y

by plugging all the values we get

ST-3.2 ± 0.83 + 3.2 = 7.23km

s, 0 + 3.09 + 0-3.09km

thus the magnitude of total displacement is

S = sqrt{s_x^2 + S_y^2}

by plugging all the values we get

S = V/(7.23)2 + (3,09)2

which gives us

S 7.86km

and total time taken during this trip is

t=t1 + t2 + t3

t = 0.1 0.21 + 0.1 0.41/27.

hence the average velocity of the entire trip is

displacement velocity- tame

7.86 0.41 0 =

19.17 ะ 19km/hr. u

Hence

the average velocity of the entire trip is

u = 19km / hr  

Option - 4 is correct answer